# Artificial Hippocampus Networks for Efficient Long-Context Modeling

### v1

@@ Abstract
Long-sequence modeling faces a fundamental trade-off between the efficiency of compressive fixed-size memory in RNN-like models and the fidelity of lossless growing memory in attention-based Transformers. Inspired by the Multi-Store Model in cognitive science, we introduce a memory framework of artificial neural networks. Our method maintains a sliding window of the Transformer’s KV cache as lossless short-term memory, while a learnable module termed Artificial Hippocampus Network (AHN) recurrently compresses out-of-window information into a fixed-size compact long-term memory. To validate this framework, we instantiate AHNs using modern RNN-like architectures, including Mamba2, DeltaNet, and GatedDeltaNet. Extensive experiments on long-context benchmarks LV-Eval and InfiniteBench demonstrate that AHN-augmented models consistently outperform sliding window baselines and achieve performance comparable or even superior to full-attention models, while substantially reducing computational and memory requirements.
@@
Figure 1: (a) Artificial Hippocampus Networks (AHNs) transform lossless memory into fixed-size compressed representations for efficient long-context modeling. Lossless memory (e.g., attention’s KV cache) stores exact input information but grows with sequence length, leading to high cost for long sequences. In contrast, compressed memory (e.g., RNNs’ hidden state) maintains a constant cache size and computational cost per input token, but inevitably loses details. In our framework, a sliding window attention maintains exact recent context as lossless short-term memory, while AHN recurrently compresses out-of-window information into a fixed-size state as compressed long-term memory. This allows the model to process long sequences efficiently, retaining both precise short-term information and a compact summary of history. (b) On the long-context benchmark LV-Eval (128k sequence length), augmenting Qwen2.5-3B-Instruct with AHNs (+0.4% parameters) reduces FLOPs by 40.5% and memory cache by 74.0%, while improving average score from 4.41 to 5.88.
@@ 1 Instruction
“Memory is the treasury and guardian of all things” [15]. Inspired by the fundamental role of memory in intelligence, researchers have long sought to model this cognitive function in artificial systems. Early efforts centered on Recurrent Neural Networks (RNNs) [23, 33, 14, 29], where sequential information is encoded by continuously updated hidden states. Over time, diverse paradigms for memory representation emerged, including key-value (KV) caches in attention mechanisms [80], external memory modules in Neural Turing Machines and Memory Networks [27, 86], and external databases for retrieval-augmented models [43]. Among these, RNN-like and attention-based models have become the most widely used, each offering distinct advantages and limitations [102, 48].
@@
When Transformers with growing lossless memory struggle for very long sequences, it is natural to revisit the RNNs’ fixed-size compressed memory, which offers constant per-token update cost regardless of context length [41, 29, 96]. This contrast highlights a fundamental trade-off between the efficiency of compressive memory and the fidelity of lossless memory. To address this problem, it is instructive to consider how the human brain maintains nearly constant volume through early and middle adulthood [19, 16, 25] while still supporting efficient processing of information across the human lifespan. The theory of Multi-Store Model of memory (MSM) in Cognitive Science and Neuroscience [4] suggests that although lossless short-term memory (or called working memory [5]) has limited capacity and duration [56, 4, 65], the hippocampus continually consolidates them into long-term cortical representations [71, 75, 3, 55, 78].
@@
To evaluate the effectiveness of AHNs, we instantiate them using Mamba2 [18], DeltaNet (DN) [70, 97] and GatedDeltaNet (GDN) [98], resulting in the AHN-Mamba2, AHN-DN and AHN-GDN. Experimental results on long-context benchmarks LV-Eval [103] and InfiniteBench [105] show that AHN-augmented models consistently outperform their sliding window counterparts, and match or even surpass full attention models while significantly reducing computational and memory cache costs. For instance, as shown in Figure 1(b), augmenting Qwen2.5-3B-Instruct [93] with AHNs (+0.4% parameters) reduces FLOPs by 40.5% and memory cache by 74.0%, while improving average score from 4.41 to 5.88 on LV-Eval (128k sequence length) [103].
@@
The contributions of this paper are twofold. First, we introduce the concept of Artificial Hippocampus Networks (AHNs), which continually transform lossless memory outside the sliding window into a compressed memory representation, enabling the model to leverage both memories for efficient long-context modeling. Second, to empirically validate the effectiveness of AHNs, we instantiate the concept into AHN-Mamba2, AHN-DN, and AHN-GDN, and train these instances using an efficient self-distillation scheme. Experimental results demonstrate that these instances substantially enhance model efficiency on long-sequence benchmarks, while achieving competitive performance compared to the full attention model. We will release the code and models to facilitate future research on the development of more AHN variants.
2 Related work
2.1 Memory in neural networks
Memory mechanisms play a crucial role in enabling neural networks to process and retain information over time, which is essential for tasks that require understanding of temporal dependencies, sequential data, or context preservation. Traditional feedforward neural networks lack the capability to maintain information across time steps, which limits their effectiveness in tasks such as language modeling, sequence prediction, and reasoning. To address this limitation, Recurrent Neural Networks (RNNs) are introduced [23, 35, 36]. RNNs maintain a hidden state that is updated at each time step, allowing information to persist across sequences. However, vanilla RNNs suffer from issues such as vanishing and exploding gradients, making it difficult to capture long-term dependencies [10]. To mitigate these problems, more advanced architectures like Long Short-Term Memory (LSTM) networks [33] and Gated Recurrent Unit (GRU) [14] are proposed. These models incorporate gating mechanisms that regulate the flow of information, enabling them to learn longer-term dependencies more effectively. Because these RNN-like models maintain a fixed-size memory and a consistent memory update cost for each input token, they are highly efficient for processing long sequences. Therefore, our AHNs are designed within the RNN paradigm to inherit this advantageous property.
2.2 Memory management
RNN-like models [23, 33, 14, 41, 64, 76, 29, 96, 97, 18, 9, 98] maintain memory through a fixed-size hidden state, regardless of input sequence length. Therefore, memory caching is not a major concern for these architectures. In contrast, Transformers store key-value (KV) pairs for every token in the input sequence, resulting in linear growth of the KV cache with sequence length. This results in significant memory consumption and presents a major challenge for processing long sequences. To mitigate this issue, various approaches have been proposed [45], including KV cache selection [26, 47, 91, 31, 107, 52, 1, 88, 79], budget allocation [12, 94, 24, 90], merging [58, 81, 83, 51], quantization [99, 74, 34, 89, 49, 72], low-rank decomposition [101, 21], external memory [62, 82], and neural architecture design [73, 2, 50, 38, 77, 100, 87, 57]. Among them, a straightforward strategy is to use a sliding window for attention [80], but this method discards KV pairs outside the window, thereby losing long-range context. Sparse Transformers [13] address this by retaining KV pairs at specific pattern positions to capture long-range dependencies, but still drop portions of the KV cache, potentially missing important information. Transformer-XL [17] introduces a segment-level recurrence mechanism by caching the last segment of hidden states as a First-In, First-Out (FIFO) memory. Compressive Transformer [68] extends this by compressing older memories into a secondary FIFO memory, but it still discards memory once the slots are full. In contrast, AHNs adopt an RNN-like paradigm that continually compresses KV pairs outside the sliding window into a lifelong compressed memory, rather than discarding them outright [48, 57, 69]. AHNs (like AHN-GDN [98]) can also dynamically control memory decay [18, 70, 97, 98]. Recent studies integrate RNNs and attention either in interleaved layers [48, 69, 18, 98, 44] or within a single layer [57, 46]. By contrast, we abstract the compression module as an AHN concept, yielding a more general memory framework. We employ a sliding-window attention mechanism, activating AHNs whenever a token leaves the window. Additionally, we introduce a simple self-distillation scheme that trains AHNs efficiently.
@@ 3.1 Preliminary
Most modern autoregressive large language models are built on Transformer architecture [80], which employs self-attention as the core mechanism for token mixing. Given an input sequence of $L$ tokens $X=(x_{1},x_{2},...,x_{L})\in\mathbb{R}^{L\times D}$, self-attention first projects the tokens into query ($Q$), key ($K$), and value ($V$) matrices via learned linear transformations:
@@ 3.2 Artificial Hippocampus Networks
Figure 2: (a) Illustration of the model augmented with Artificial Hippocampus Networks (AHNs). In this example, the sliding window length is 3. When the input sequence length is less than or equal to the window length, the model operates identically to a standard Transformer. For longer sequences, AHNs continually compress the token outside the window into a compact memory representation. The model then utilizes both the lossless information within window, and the compressed memory to generate the next token. (b) Self-distillation training framework of AHNs based on an open-weight LLM. During training, the base LLM’s weights are frozen, and only the AHNs’ parameters are trained.
@@ 3.3 Instantiation
As discussed above, AHNs can be instantiated using RNN-like architectures. In our experiments, we focus on modern linear recurrent models for their efficient parallel training. Specifically, we utilize three architectures including Mamba2 [18], DeltaNet (DN) [70, 97], and its enhanced version, GatedDeltaNet (GDN) [96], to instantiate AHNs into AHN-Mamba2, AHN-DN and AHN-GDN, respectively. Below, we present the implementation of AHN-GDN as a representative example, and the other two AHN instances are described in Appendix 6. Specifically, AHN-GDN updates memory via the gated delta rule [70, 97, 96]:
@@
Unlike GatedDeltaNet [98], which compresses all past tokens, AHN-GDN only compresses tokens outside the sliding window. For each position $t$, the query $q_{t}$ derived from $x_{t}$ is used to access the compressed memory $h_{t-W}$. The output is further modulated by a gate function $\gamma(x_{t})$ and then is transformed by a linear projection:
@@
Different from GatedDeltaNet [96], the output of $\gamma(x_{t})$ is a scalar shared across head channels, and the output linear is grouped by heads [42, 39] with learnable weight $W_{o}\in\mathbb{R}^{H\times H}$ ($H$ denotes head dimension). Finally, we simply sum the outputs from AHN and the attention mechanism:
@@ 3.4 Training framework
While an AHN-augmented model can be trained from scratch, we adopt a more computationally efficient approach using self-distillation [32, 106, 104]. This allows us to leverage powerful pre-trained models. Our training framework uses an open-weight LLM (e.g., Qwen [93]) as the teacher model, with its output probability denoted as $p^{\prime}$. The student model is the same LLM, but we modify its attention mechanism to operate over a limited receptive field of a sliding window at every layer. These window attention layers are then augmented with AHNs. The student’s output probability is denoted as $p$. We train the student to mimic the teacher’s output distribution by minimizing the Kullback-Leibler (KL) divergence:
$l=\text{KL}(p^{\prime}||p).$ (8)
To maximize efficiency, the base model’s weights are frozen during training, and only the AHN parameters are optimized. This framework is illustrated in Figure 2(b).
@@ 4.1 Setups
Implementation details. We implement all AHN instances in PyTorch [63], building on LLaMA-Factory [108] and Flash Linear Attention [95]. During training, we freeze the base LLM and train the newly initialized AHN module using a self-distillation loss, as illustrated in Figure 2(b). To ensure the AHN module learns a generalizable compression strategy, we randomize the starting position of the AHN modules and also the sliding window size. For optimization, we use the AdamW [53] optimizer with a learning rate of 1e-4, which is warmed up linearly over the first 10% of steps and then cosine decayed. All models are trained for one epoch on the ChatQA2 dataset, using a global batch size of 128.
@@
Sinks + SWA 0% 48.0% 65.2% 25.6% 9.52 4.76 14.09 5.34 10.66 15.66 13.16
CT-Max 0% 48.5% 65.6% 26.0% 8.35 4.02 12.34 4.82 10.56 15.45 13.00
CT-Average 0% 48.5% 65.6% 26.0% 9.48 4.86 13.78 5.28 10.63 15.99 13.31
AHN-Mamba2 0.2% 48.1% 65.4% 26.0% 12.57 5.54 14.13 6.21 11.36 17.06 14.21
AHN-DN 0.2% 48.1% 65.4% 26.0% 11.97 5.67 16.52 6.82 12.86 20.10 16.48
AHN-GDN 0.3% 48.1% 65.4% 26.0% 12.69 4.71 15.30 6.54 13.37 20.48 16.93
@@ 4.4 Ablation study
Having demonstrated the effectiveness of AHN-augmented models, we now conduct an ablation study to analyze the impact of our two design choices: the training objective and the use of randomization. For these experiments, we use AHN-GDN (Qwen2.5-7B-Instruct) as the starting point.
@@
Figure 4: AHN modules demonstrate strong context generalization capacity on LongBench.
Table 4: Ablation of AHN training design choices. We ablate two factors: (1) the training objective, comparing self-distillation (KL loss) with next-token prediction (no full-attention teacher model, CE loss), and (2) randomized versus fixed sliding window configurations. All experiments are based on Qwen2.5-7B-Instruct with AHN-GDN.
@@
Self-distillation (KL loss) 1024 (fixed) 38.53
Next-token prediction (CE loss) Random size 39.59
Self-distillation (KL loss) Random size 40.59
Randomization vs. fixed windows. We train AHNs with randomized sliding window sizes to encourage a general compressive module that adapts to varying look-ahead contexts. By comparison, models trained with fixed windows tend to overfit to the specific configuration and fail to generalize to unseen context lengths. This suggests that the design of compression priors and training schedules deserves further investigation. To evaluate context generalization, we fix attention sinks to 128 tokens and test AHN-augmented models with sliding window sizes from 896 to 8064. As shown in Figure 4, AHN-augmented models maintain strong performance across all tested configurations.

### v2

@@ Abstract
[…]
Abstract: Long-sequence modeling faces a fundamental trade-off between the efficiency of compressive fixed-size memory in RNN-like models and the fidelity of lossless growing memory in attention-based Transformers. Inspired by the Multi-Store Model in cognitive science, we introduce a memory framework of artificial neural networks. Our method maintains a sliding window of the Transformer’s KV cache as lossless short-term memory, while a learnable module termed Artificial Hippocampus Network (AHN) recurrently compresses out-of-window information into a fixed-size compact long-term memory. To validate this framework, we instantiate AHNs using modern RNN-like architectures, including Mamba2, DeltaNet, and GatedDeltaNet to augment open-weight LLMs. We also propose an efficient self-distillation training method where the base model’s all parameters are frozen and only the parameters from AHNs are optimized. For inference, our method sets a default large sliding window size of 32k for attention, and AHNs activate only when the sequence length exceeds the 32k window, addressing the quadratic-complexity issue of attention that emerges at that scale. Extensive experiments on long-context benchmarks LV-Eval and InfiniteBench demonstrate that AHN-augmented models consistently outperform sliding window baselines and achieve performance comparable or even superior to full-attention models, while substantially reducing computational and memory requirements.
@@
Figure 1: (a) Artificial Hippocampus Networks (AHNs) transform lossless memory into fixed-size compressed representations for efficient long-context modeling. Lossless memory (e.g., attention’s KV cache) stores exact input information but grows with sequence length, leading to high cost for long sequences. In contrast, compressed memory (e.g., RNNs’ hidden state) maintains a constant cache size and computational cost per input token, but inevitably loses details. In our framework, a sliding window attention maintains exact recent context as lossless short-term memory, while AHN recurrently compresses out-of-window information into a fixed-size state as compressed long-term memory. This allows the model to process long sequences efficiently, retaining both precise short-term information and a compact summary of history. See Figure 2 for more details. (b) On the long-context benchmark LV-Eval (128k sequence length), augmenting Qwen2.5-3B-Instruct with AHNs (+0.4% parameters) reduces FLOPs by 40.5% and memory cache by 74.0%, while improving average score from 4.41 to 5.88.
@@ 1 Instruction
“Memory is the treasury and guardian of all things” [16]. Inspired by the fundamental role of memory in intelligence, researchers have long sought to model this cognitive function in artificial systems. Early efforts centered on Recurrent Neural Networks (RNNs) [92, 43, 25, 35, 15, 37, 38], where sequential information is encoded by continuously updated hidden states. Over time, diverse paradigms for memory representation emerged, including key-value (KV) caches in attention mechanisms [85], external memory modules in Neural Turing Machines and Memory Networks [29, 93], and external databases for retrieval-augmented models [47]. Among these, RNN-like and attention-based models have become the most widely used, each offering distinct advantages and limitations [109, 52].
@@
When Transformers with growing lossless memory struggle for very long sequences, it is natural to revisit the RNNs’ fixed-size compressed memory, which offers constant per-token update cost regardless of context length [45, 31, 103]. This contrast highlights a fundamental trade-off between the efficiency of compressive memory and the fidelity of lossless memory. To address this problem, it is instructive to consider how the human brain maintains nearly constant volume through early and middle adulthood [20, 17, 27] while still supporting efficient processing of information across the human lifespan. The theory of Multi-Store Model of memory (MSM) in Cognitive Science and Neuroscience [4] suggests that although lossless short-term memory (or called working memory [5]) has limited capacity and duration [61, 4, 70], the hippocampus continually consolidates them into long-term cortical representations [76, 80, 3, 60, 24, 83].
@@
To evaluate the effectiveness of AHNs, we instantiate them using Mamba2 [19], DeltaNet (DN) [75, 104] and GatedDeltaNet (GDN) [105], resulting in the AHN-Mamba2, AHN-DN and AHN-GDN. We introduce an efficient self-distillation training scheme in which the teacher model is an open-weight attention-based model (e.g., Qwen), and the student model shares the teacher’s parameters but with token mixer of window attention and AHN. We employ a KL divergence loss, optimizing only the AHN parameters while freezing all remaining parameters, as shown in Figure 2(b). The models on trained on ChatQA 2 [99] with 1B tokens, sample sequence length up to 24k, and random sliding window size up to 8k, which only cost $\sim 10$ hours on 32 A100 GPUs to train AHNs to augment 7B model. Notably, for inference, we set a default sliding-window attention size of 32k, which is substantially larger than those used in prior attention–RNN hybrid methods (e.g., 64 in [112, 41]) AHNs activate only when the sequence length exceeds the 32k window, addressing the quadratic-complexity issue of attention that emerges at that scale.
@@
Experimental results on long-context benchmarks LV-Eval [110] and InfiniteBench [113] show that AHN-augmented models consistently outperform their sliding window counterparts, and match or even surpass full attention models while significantly reducing computational and memory cache costs. For instance, as shown in Figure 1(b), augmenting Qwen2.5-3B-Instruct [100] with AHNs (+0.4% parameters) reduces FLOPs by 40.5% and memory cache by 74.0%, while improving average score from 4.41 to 5.88 on LV-Eval (128k sequence length) [110].
@@
The contributions of this paper are twofold. First, we introduce the concept of Artificial Hippocampus Networks (AHNs), which continually transform lossless memory outside the sliding window into a compressed memory representation, enabling the model to leverage both memories for efficient long-context modeling. Second, to empirically validate the effectiveness of AHNs, we instantiate the concept into AHN-Mamba2, AHN-DN, and AHN-GDN, and train these instances using an efficient self-distillation scheme. Experimental results demonstrate that these instances substantially enhance model efficiency on long-sequence benchmarks, while achieving competitive performance compared to the full attention model.
@@ 2.1 Preliminary
Most modern autoregressive large language models are built on Transformer architecture [85], which employs self-attention as the core mechanism for token mixing. Given an input sequence of $L$ tokens $X=(x_{1},x_{2},...,x_{L})\in\mathbb{R}^{L\times D}$ ($D$ is the hidden dimension), self-attention first projects the tokens into query ($Q$), key ($K$), and value ($V$) matrices via learned linear transformations:
@@ 2.2 Artificial Hippocampus Networks
Figure 2: (a) Illustration of the model augmented with Artificial Hippocampus Networks (AHNs). In this illustrative example, we set the sliding window length to 3 for clarity. For model inference in our experiments, the default window length is 32k. When the input sequence length is less than or equal to the window length, the model operates identically to a standard Transformer. For longer sequences, AHNs continually compress the token outside the window into a compact memory representation. The model then utilizes both the lossless information within window, and the compressed memory to generate the next token. (b) Self-distillation training framework of AHNs based on an open-weight LLM. During training, the base LLM’s weights are frozen, and only the AHNs’ parameters are trained.
@@ 2.3 Instantiation
As discussed above, AHNs can be instantiated using RNN-like architectures. In our experiments, we focus on modern linear recurrent models for their efficient parallel training. Specifically, we utilize three architectures including Mamba2 [19], DeltaNet (DN) [75, 104], and its enhanced version, GatedDeltaNet (GDN) [103], to instantiate AHNs into AHN-Mamba2, AHN-DN and AHN-GDN, respectively. Below, we present the implementation of AHN-GDN for each head as a representative example, and the other two AHN instances are described in Appendix 6. Specifically, AHN-GDN updates memory via the gated delta rule [75, 104, 103]:
@@
where learnable parameters for per head are $W_{\alpha}\in\mathbb{R}^{D\times 1}$ in $\alpha(\cdot)$ and $W_{\beta}\in\mathbb{R}^{D\times 1}$ in $\beta(\cdot)$. Unlike GatedDeltaNet [105], which compresses all past tokens, AHN-GDN only compresses tokens outside the sliding window. For each position $t$, the query $q_{t}$ derived from $x_{t}$ is used to access the compressed memory $h_{t-W}$. Note that AHNs do not introduce separate QKV projection layers. Instead, they directly transform the lossless memory (i.e., the KV cache) from attention into a fixed-size compact memory. The compressed memory $h_{t-W}$ is further modulated by a gate function $\gamma(x_{t})$ and then is transformed by a linear projection to generate output:
@@
Different from GatedDeltaNet [105], the output of $\gamma(x_{t})=x_{t}W_{\gamma}$ is a scalar for each head with learnable parameter $W_{\gamma}\in\mathbb{R}^{D\times 1}$, and the output linear is grouped by heads [46, 42] with learnable weight $W_{o}\in\mathbb{R}^{H\times H}$ ($H$ denotes head dimension). Finally, we simply sum the outputs from AHN and the attention mechanism:
@@ 2.4 Training framework
While an AHN-augmented model can be trained from scratch, we adopt a more computationally efficient approach using self-distillation [34, 114, 111]. This allows us to leverage powerful pre-trained models. Our training framework uses an open-weight LLM (e.g., Qwen [100]) as the teacher model, with its output probability denoted as $p^{\prime}$. The student model is the same LLM, but we modify its attention mechanism to operate over a limited receptive field of a sliding window at every layer. These window attention layers are then augmented with AHNs. The student’s output probability is denoted as $p$. We train the student to mimic the teacher’s output distribution by minimizing the Kullback-Leibler (KL) divergence: $l=\text{KL}(p^{\prime}||p)$. To maximize efficiency, the base model’s weights are frozen during training, and only the AHN parameters are optimized. Taking AHN-GDN as an example, only the parameters involved in Equations 5 and 6 are learnable. For each attention head, these trainable parameters consist of the gating weights $W_{\alpha}\in\mathbb{R}^{D\times 1}$ in $\alpha(\cdot)$, $W_{\beta}\in\mathbb{R}^{D\times 1}$ in $\beta(\cdot)$, $W_{\gamma}\in\mathbb{R}^{D\times 1}$ in $\gamma(\cdot)$ as well as the output projection $W_{o}\in\mathbb{R}^{H\times H}$. Here, $D$ and $H$ denote the hidden dimension and the head dimension, respectively. With $N_{\text{q}}$ attention heads, the model contains $N_{\text{q}}$ such sets of parameters, amounting to only $\sim 0.4\%$ relative to the frozen base model’s parameters. The framework is illustrated in Figure 2(b).
@@ 3.1 Setups
Implementation details. We implement all AHN instances in PyTorch [68], building on LLaMA-Factory [117] and Flash Linear Attention [102]. During training, we freeze the entire base LLM and only train the newly added AHN module (only $\sim 0.4\%$ parameters relative to base LLM) using self-distillation, as illustrated in Figure 2(b). To ensure the AHN module learns a generalizable compression strategy, we randomize the starting position (the number of attention sinks) of the AHN modules and also the sliding window size. Specifically, we use a maximum sequence length of 24k tokens during training. For each example, the attention-sink size is uniformly sampled from [0, 32, 64, 128, 512, 2048, 4096] after removing any candidates larger than half of the sequence length. The total token number of lossless memory (attention sinks + sliding window) is uniformly sampled from [32, 64, 128, 256, 512, 1024, 2048, 4096, 8192] after filtering out values smaller than one-eighth of the sequence length. For optimization, we use the AdamW [58] optimizer with a learning rate of 1e-4, which is warmed up linearly over the first 10% of steps and then cosine decayed. All models are trained for one epoch on the ChatQA2 dataset, consisting of 1B tokens, using a global batch size of 128 for a total of 740 update steps. Training AHNs for a 7B base model requires only $\sim$10 hours on 32 A100 GPUs.
@@
Sinks + SWA 0% 48.0% 65.5% 25.6% 9.52 4.76 14.09 5.34 10.66 15.66 13.16
CT-Max 0% 48.5% 65.8% 26.0% 8.35 4.02 12.34 4.82 10.56 15.45 13.00
CT-Average 0% 48.5% 65.8% 26.0% 9.48 4.86 13.78 5.28 10.63 15.99 13.31
AHN-Mamba2 0.2% 48.2% 65.6% 26.0% 12.57 5.54 14.13 6.21 11.36 17.06 14.21
AHN-DN 0.2% 48.2% 65.6% 26.0% 11.97 5.67 16.52 6.82 12.86 20.10 16.48
AHN-GDN 0.3% 48.2% 65.6% 26.0% 12.69 4.71 15.30 6.54 13.37 20.48 16.93
@@ 3.4 Ablation study
Having demonstrated the effectiveness of AHN-augmented models, we now conduct an ablation study to analyze the impact of our three design choices: the training objective, randomization of training window size, and the inference window size. For these experiments, we use AHN-GDN (Qwen2.5-7B-Instruct) as the starting point.
@@
Figure 4: The AHN-augmented model consistently outperforms the sliding-window attention (SWA) baseline across all window sizes on both LV-Eval and InfiniteBench (128k sequence length).
Table 4: Ablation of AHN training design choices based on Qwen2.5-7B-Instruct: (1) the training objective (2) randomized vs. fixed window.
@@
Self-distill (KL loss) 1024 (fixed) 38.53
Next-token pred. (CE loss) Random size 39.59
Self-distil. (KL loss) Random size 40.59
Training window size: randomized vs. fixed. We train AHNs using randomized sliding-window sizes to encourage a more general and robust compressive module that can adapt to varying context lengths. In contrast, models trained with a fixed window often overfit to that specific configuration and fail to generalize to unseen context lengths, leading to noticeable performance degradation, as shown in Table 4.
Inference window size. To further evaluate context length generalization, we fix the attention sinks to 128 tokens and test AHN-augmented models with sliding window sizes ranging from 1k to 96k on the 128k-context subsets of both LV-Eval and InfiniteBench. As shown in Figure 4, the AHN-augmented model maintains competitive performance across all window configurations, and consistently outperforms sliding window attention (SWA). Notably, as the inference window size increases from 1k to 16k, the performance improves steadily, highlighting the importance of a large attention window for extra-long context tasks. Beyond this range, however, we observe a noticeable performance drop, after 64k on LV-Eval and 96k on InfiniteBench, which may be attributed to the attention-dilution effect, where the the attention distribution becomes overly diffuse when the number of keys grows very large, weakening the model’s ability to focus on relevant information. Balancing performance and computational cost, we therefore adopt a 32k sliding window as the default configuration for inference. Additional LongBench results are provided in the appendix.